Calculate the pH of a Sr(OH)2 solution prepared by dissolving 0.3211g of Sr(OH)2 and making up to 250.00mL solution
1
Expert's answer
2021-10-17T11:17:29-0400

Moles of Sr(OH)2 in 250ml.


"\frac{mass}{molar\space mass}=\frac{0.3211}{121.63}=0.0026399736"

We can use the following relation to find molarity.


0.0026399736 moles "\equiv" 250 ml


For 1000 ml we have

"\frac{1000}{250}×0.0026399736=0.01056 \space moles"

pOH=-log [Sr (OH)2]

pH for bases is essentially 14-pOH.

"pH=14-(-log\space0.01056)=12.024"